# MuJo: Multimodal Joint Feature Space Learning for Human Activity Recognition

### v2

@@ Abstract
Human Activity Recognition (HAR) is a longstanding problem in AI with applications in a broad range of areas, including healthcare, sports and fitness, security, and more. The performance of HAR in real-world settings is strongly dependent on the type and quality of the input signal that can be acquired. Given an unobstructed, high-quality camera view of a scene, computer vision systems, in particular in conjunction with foundation models, can today fairly reliably distinguish complex activities. On the other hand, recognition using modalities such as wearable sensors (which are often more broadly available, e.g., in mobile phones and smartwatches) is a more difficult problem, as the signals often contain less information and labeled training data is more difficult to acquire. To alleviate the need for labeled data, we introduce our comprehensive Fitness Multimodal Activity Dataset (FiMAD) in this work, which can be used with the proposed pre-training method MuJo (Multimodal Joint Feature Space Learning) to enhance HAR performance across various modalities. FiMAD was created using YouTube fitness videos and contains parallel video, language, pose, and simulated IMU sensor data. MuJo utilizes this dataset to learn a joint feature space for these modalities. We show that classifiers pre-trained on FiMAD can increase the performance on real HAR datasets such as MM-Fit, MyoGym, MotionSense, and MHEALTH. For instance, on MM-Fit, we achieve an Macro $F_{1}$-Score of up to 0.855 when fine-tuning on only 2% of the training data and 0.942 when utilizing the full training set for classification tasks. We have compared our approach to other self-supervised ones and showed that, unlike them, ours can consistently improve on the baseline network performance as well as provide a better data-efficiency.
{NoHyper}00footnotetext: Equal contribution.
@@ 1 Introduction
Perceiving and interpreting human activities is a core functionality of intelligent systems with applications in a broad range of areas, including healthcare Tan et al. (2021), sports and fitness Host and Ivašić-Kos (2022); Nadeem et al. (2020), and security Sunil et al. (2021), among others. A basic capability needed to facilitate such functionality is the identification of concrete human actions from the system’s sensory input, referred to as Human Activity Recognition (HAR). In general, HAR is a difficult problem, especially when applied to uncontrolled real-life environments, where it must contend with the variability of human actions and the open-ended variations in environmental conditions. In such settings, the performance of HAR systems is strongly dependent on the available input modalities. Recent advances in Computer Vision, particularly with Large Vision-Language Models (LVLMs) Radford et al. (2021); Achiam et al. (2023), have enabled impressive performance when given sufficiently high-quality images. In many cases, merely feeding images to publicly available tools such as GPT-4 can result in accurate description of scenes involving complex and subtle activities. However, in real-life settings, such images are not always available; either due to a lack of cameras or because of privacy concerns. Much more broadly available input modalities are wearable sensors, in particular, Inertial Measurement Units (IMUs), which are ubiquitous in smartphones, smartwatches or even earbuds, all of which people carry with them on a daily basis. Unfortunately, achieving acceptable recognition performance with such sensors is a much more challenging problem due to two factors, the first one being information ambiguity. An IMU (a motion sensor) in a smartwatch on the wrist or in a smartphone in a pocket contains much less detailed information about the user’s activity than a high-quality picture. This problem is heightened by the second factor, lack of training data. The success of LVLMs is built on the online availability of massive amounts of text/image data pairs, which can be used for training aligned representations. Although large amounts of unlabeled IMU data are increasingly becoming available Chan et al. (2024), labeled training data remains scarce.
The work presented in this paper, which aims to address these issues, is founded on three key observations:
@@
    While the input modalities availability during inference—the time when the system is utilized in the real world—may be constrained, the input modalities during training can be selected with much greater flexibility.
@@
    Significant progress has recently been made in generating synthetic data for a variety of sensors from videos. For example, IMUTube Kwon et al. (2021) generates virtual gyroscope and accelerometer data from videos, while pose extraction methods like OpenPose Cao et al. (2017) and VideoPose3D Pavllo et al. (2019) are used to generate skeletal data of human activities. This enables the creation of synchronized video/sensor training data and through the alignment of video with text, it facilitates the synchronization of sensor data with textual descriptions.
Based on the above considerations, we propose a Multimodal Joint Feature Space Learning method (MuJo) that learns a joint representation between videos, video-derived poses, virtual IMU data generated from videos, and textual descriptions by employing pair-wise contrastive learning, as depicted in Figure 1. MuJo aims to utilize the complementary information of each modality to improve the performance for both unimodal and multimodal inference. In particular, we show that our method dramatically reduces the need for labeled training data for downstream tasks. For that, we provide extensive evaluation on four publicly available HAR datasets, namely MM-Fit Strömbäck et al. (2020), MyoGym Koskimäki et al. (2017), MotionSense Malekzadeh et al. (2018), and MHEALTH Banos et al. (2015). For instance, on the MM-Fit dataset in the multimodal case (encompassing all of the aforementioned modalities as input), our method improves the recognition rate (Macro $F_{1}$-Score) from 0.897 to 0.942 when using all available training data. Even when utilizing only 2% of the training data, it enhances the recognition rate from 0.790 to 0.855.
@@
    Development of the Fitness Multimodal Activity Dataset (FiMAD) [^1] Available after publication at link., a comprehensive dataset containing parallel video, language (including video clip labels, instructor utterances, and detailed exercise descriptions generated by GPT-3.5 Turbo), pose data, and simulated IMU data points.
@@
    Pe-training of MuJo, a multimodal joint space model, using the FiMAD dataset. MuJo integrates four different modalities: text, video, pose, and virtual sensor data, leveraging contrastive learning inspired by CLIP Radford et al. (2021).
@@
    Evaluation of the utility and generalizability of MuJo for real-world accelerometer data on the MM-Fit, MyoGym, MotionSense, and MHEALTH datasets. On the MM-Fit dataset, we additionally conducted experiments with pose, video, and multimodal inputs. The benefits of pre-training MuJo on FiMAD, especially when limited data is available for the target task, are analyzed. Furthermore, we analyze the zero-shot performance of our method and compare unimodal versus multimodal training on all four datasets.
@@ 2 Related Work
Unimodal and multimodal activity recognition.
@@
Unimodal activity recognition has been extensively studied in HAR. Various approaches have been proposed to address this task, and they can be broadly categorized into two groups: hand-crafted feature-based and deep learning-based methods. Hand-crafted feature-based methods rely on extracting low-level features from the input data and designing a classifier to recognize activities. For instance, in accelerometer-based activity recognition Kwapisz et al. (2011), features such as mean, standard deviation, and energy of the acceleration signal are often used to capture the characteristics of different activities. Popular classifiers used in this approach include decision trees, support vector machines, and k-nearest neighbors. Deep learning-based methods, on the other hand, use neural networks to learn features and recognize activities from raw input data. Convolutional neural networks (CNNs) and recurrent neural networks (RNNs) are two popular types of neural networks used in this approach. CNNs are commonly used for image-based activity recognition Strömbäck et al. (2020), being combined with RNNs in some cases Ordóñez and Roggen (2016). Despite the success of unimodal activity recognition, it is often limited by the fact that a single modality of input data may only capture some of the necessary information about the activity, leading to lower recognition accuracy. Therefore, multimodal activity recognition, which combines information from multiple sources, has gained increasing attention in recent years.
@@
Multimodal activity recognition, on the other hand, recognizes human activities using information from multiple modalities, such as IMU data, audio, and video Mekruksavanich and Jitpattanakul (2022); Ijaz et al. (2022). These features can then be combined using fusion methods such as early fusion or late fusion Duhme et al. (2022). Some studies have also focused on domain adaptation, where the goal is to recognize activities in a new environment using data from a different environment Plananamente et al. (2022). In such cases, transfer learning techniques can adapt the model to the new environment using a small amount of labeled data. Overall, multimodal activity recognition has shown promising results, with improved accuracy compared to unimodal approaches. However, challenges in data fusion, feature extraction, and model selection still need to be addressed to improve the performance of multimodal activity recognition systems Islam et al. (2022).
Representation learning with proxy tasks.
@@
Different modalities, e.g., natural language, visual inputs, or sensor signals, often complement a common concept. A key challenge of multimodal feature representation is exploring efficient methods for multimodal fusion with the given modalities to avoid missing modal features. We can differentiate the representation learning methods in two aspects: learning objectives and fusing the learned features in a joint space. In learning objectives, various approaches have been proposed to learn multimodal representations by formulating proxy tasks such as reconstruction, completion Seo et al. (2021); Sun et al. (2019), matching/alignment Li et al. (2020); Miech et al. (2020), and ordering tasks Li et al. (2020). Generating one feature vector with multimodal inputs for a common concept is a crucial problem of multimodal representation learning. Simple methods, such as concatenating learned uni-modal feature vectors, have been extensively explored to enhance the performance of downstream tasks Hu et al. (2018); Strömbäck et al. (2020). These methods involve combining feature vectors from a single modality, such as text or image, to create more informative representations that can be used for various tasks. More advanced methods have also been proposed, including a tensor fusion network Zadeh et al. (2017) and a transformer-based co-attention mechanism for efficient cross-modality learning Lu et al. (2019).
@@
Proxy task were also developed for the field of Human Activity Recognition with weareable devices including IMUs, where unlabelled IMU data is used to pre-train better representations. Those approaches include Multitask Saeed et al. (2019), Auto-encoder Haresamudram et al. (2019) and SimClr Tang et al. (2021). Multi-task Saeed et al. (2019) proposes auxiliary tasks of binary classification for extracting useful features for the downstream task. A multi-task temporal convolutional neural network (CNN) is trained to classify whether the input signal was transformed or not for each of a set of transformations such as scaling, rotating the signal, etc. The authors utilize a large unlabelled accelerometer dataset for self-supervised learning and transfer learned knowledge to a small, labelled activity recognition dataset. Auto-encoder Haresamudram et al. (2019) is the application of auto-encoders where an encoder network reduces the dimensionality of the data while a decoder one brings the representation back to its original dimensions, with the whole system learning to reconstruct the original signal. SimClr Tang et al. (2021) learns general representations from a dataset where a model is trained to match different views instances created from the same signal by contrasting them. The positive pair is obtained by the transformation of the input window, while the negative pair is formed by the rest input windows of the same batch. Then, the contrastive loss pulls the positive pair together while pulling the negative pair far away from each other in the representation space.
Representation via contrastive learning.
@@
Contrastive learning is as an efficient self-supervised framework applied across multiple domains, which learns similar/dissimilar representations from data that are organized into similar and dissimilar pairs. Recently, Radford et al. presented a novel approach, known as CLIP Radford et al. (2021), to jointly represent images and text descriptions by leveraging contrastive loss with the similarity of image-text pairs. CLIP learns two feature spaces with two encoders, i.e., image and text encoders, then projects the learned features into a shared latent space with contrastive loss. Contrastive learning has been extended to other modalities as well. For example, video-text data has attracted interest in studies focused on the shared space of video-text features Xu et al. (2021); Xue et al. (2022) and text-to-video generation Singer et al. (2022). MotionCLIP Tevet et al. (2022) trained an encoder to find the appropriate embedding of an input sequence in CLIP space and a decoder to generate the most fitting motion for a given CLIP space. Following this, IMU2CLIP Moon et al. (2022) was proposed to learn a joint space that aligns IMU motion sensor recordings with video and text by projecting them into a shared representation space.
Studies in action recognition Li et al. (2020); Wang et al. (2021); Sun et al. (2022); Morshed et al. (2023) have also significantly enhanced the field of fitness activity recognition, despite traditionally focusing on a broad spectrum of actions such as “cutting the potato” and “cleaning the drawer”. Certain investigations in action recognition also endeavor to acquire knowledge of the feature space through contrastive learning, as exemplified by initiatives like HERO Li et al. (2020) and ActionCLIP Wang et al. (2021).
@@
The achievements discussed in the aforementioned studies have shown that contrastive learning has a significant impact on extracting a joint space with similar and dissimilar pairs. Consequently, the contrastive learning methodology employed in our study aims to acquire knowledge of the feature space pre-training to fitness activities. Contrary to previous work, we extend this line of work on contrastive learning to a unique multimodal setting that utilizes more than two modalities, such as video, IMU sensors, poses, and text, within a multitask training manner in an end-to-end training paradigm similar to ImageBind Girdhar et al. (2023). Thus, our model and data provide a cross-modal classification opportunity.

### v3

@@ Abstract
Human activity recognition (HAR) is a long-standing problem in artificial intelligence with applications in a broad range of areas, including healthcare, sports and fitness, security, and more. The performance of HAR in real-world settings is strongly dependent on the type and quality of the input signal that can be acquired. Given an unobstructed, high-quality camera view of a scene, computer vision systems, in particular in conjunction with foundation models, can today fairly reliably distinguish complex activities. On the other hand, recognition using modalities such as wearable sensors (which are often more broadly available, e.g., in mobile phones and smartwatches) is a more difficult problem, as the signals often contain less information and labeled training data is more difficult to acquire. To alleviate the need for labeled data, we introduce our comprehensive Fitness Multimodal Activity Dataset (FiMAD) in this work, which can be used with the proposed pre-training method MuJo (Multimodal Joint Feature Space Learning) to enhance HAR performance across various modalities. FiMAD was created using YouTube fitness videos and contains parallel video, language, pose, and simulated IMU sensor data. MuJo utilizes this dataset to learn a joint feature space for these modalities. We show that classifiers pre-trained on FiMAD can increase the performance on real HAR datasets such as MM-Fit, MyoGym, MotionSense, and MHEALTH. For instance, on MM-Fit, we achieve a Macro $F_{1}$-Score of up to 0.855 when fine-tuning on only 2% of the training data and 0.942 when utilizing the complete training set for classification tasks. We compare our approach with other self-supervised ones and show that, unlike them, ours consistently improves compared to the baseline network performance while also providing better data efficiency.
{NoHyper}00footnotetext: This research was supported by the XAINES project, funded by the German Federal Ministry of Education and Research under grant agreement No 01IW20005, and by the SustainML project, funded by the European Union’s Horizon Europe research and innovation programme under grant agreement No 101070408.
@@ 1 Introduction
Perceiving and interpreting human activities is a core functionality of intelligent systems with applications in a wide range of areas, including healthcare (Tan et al., 2021), sports and fitness (Host and Ivašić-Kos, 2022; Nadeem et al., 2020), and security (Sunil et al., 2021), among others. A basic capability needed to facilitate such functionality is the identification of concrete human actions from the system’s sensory input, referred to as human activity recognition (HAR). In general, HAR is a difficult problem, especially when applied to uncontrolled real-life environments, where it must contend with the variability of human actions and the open-ended variations in environmental conditions. In such settings, the performance of HAR systems is strongly dependent on the available input modalities. Recent advances in computer vision, particularly with large vision-language models (LVLM) (Radford et al., 2021; Achiam et al., 2023), have enabled impressive performance when provided with sufficiently high-quality images. In many cases, simply feeding images to publicly available tools such as GPT-4 can result in accurate descriptions of scenes that involve complex and subtle activities. However, in real-life settings, such images are not always available; either due to a lack of cameras or because of privacy concerns. Much more broadly available input modalities are wearable sensors, in particular, inertial measurement units (IMUs), which are ubiquitous in smartphones, smartwatches, or even earbuds, all of which people carry with them on a daily basis. In contrast to cameras, such sensors are more privacy-preserving, since they collect only data from the subject who wears them. Moreover, they are cost- and energy-efficient and, for instance, in the case of smartwatches and smartphones, convenient to wear. Unfortunately, achieving acceptable recognition performance with such sensors is a much more challenging problem due to two factors, the first being information ambiguity. An IMU (a motion sensor) in a smartwatch on the wrist or in a smartphone in a pocket contains much less detailed information about the user’s activity than a high-quality picture. This problem is heightened by the second factor, the lack of training data. IMU-based HAR datasets are very small compared to vision HAR ones. For example, the total length of all sensor-based HAR datasets in the popular UCI machine learning repository (Kelly et al., 2025) is less than 63 hours (Fortes Rey et al., 2021), which pales in comparison to a single video HAR dataset such as Kinetics-700 with around 1800 hours (Carreira et al., 2019). Moreover, the massive amount of text/image and text/video data pairs available online led to the success of LVLMs, which can be used for training aligned representations that capture the semantic similarity of multiple modalities. On the other hand, although large amounts of unlabeled IMU data are increasingly becoming available (Chan et al., 2024), labeled training data remain scarce. The work presented in this paper, which aims to address these issues, is based on three key observations:
@@
    Although the availability of input modalities during inference, the time when the system is utilized in the real world, may be constrained, the input modalities during training can be selected with much greater flexibility.
@@
    Significant progress has recently been made in generating synthetic data for a variety of sensors from videos. For example, IMUTube (Kwon et al., 2021a) generates virtual gyroscope and accelerometer data from videos, while pose extraction methods such as OpenPose (Cao et al., 2017) and VideoPose3D (Pavllo et al., 2019) are used to generate skeletal data of human activities. This enables the creation of synchronized video/sensor training data, and through the alignment of video with text, it facilitates the synchronization of sensor data with textual descriptions.
Based on the above considerations, we propose a Multimodal Joint Feature Space Learning method (MuJo) that learns a joint representation between videos, video-derived poses, virtual IMU data generated from videos, and textual descriptions by employing pair-wise contrastive learning, as depicted in Figure 1. MuJo aims to utilize the complementary information of each modality to improve the performance for both unimodal and multimodal inference. In particular, we show that our method dramatically reduces the need for labeled training data for downstream tasks. For that, we provide an extensive evaluation on four publicly available HAR datasets, namely MM-Fit (Strömbäck et al., 2020), MyoGym (Koskimäki et al., 2017), MotionSense (Malekzadeh et al., 2018), and MHEALTH (Banos et al., 2015). For instance, on the MM-Fit dataset in the multimodal case (encompassing all the aforementioned modalities as input), our method improves the recognition rate (Macro $F_{1}$-Score) from 0.897 to 0.942 when using all available training data. Even when utilizing only 2% of the training data, it enhances the recognition rate from 0.790 to 0.855.
@@
    Development of the Fitness Multimodal Activity Dataset (FiMAD) [^1] https://github.com/R-tree/MuJo.git, a comprehensive dataset containing parallel video, language (including video clip labels, instructor utterances, and detailed exercise descriptions generated by GPT-3.5 Turbo), pose data, and simulated IMU data points.
@@
    Pre-training of MuJo1, a multimodal joint space model, using the FiMAD dataset. MuJo integrates four different modalities: text, video, pose, and virtual sensor data, leveraging contrastive learning inspired by CLIP (Radford et al., 2021).
@@
    Evaluation of the utility and generalizability of MuJo for real-world accelerometer data on the MM-Fit, MyoGym, MotionSense, and MHEALTH datasets. On the MM-Fit dataset, we additionally conducted experiments with pose, video, and multimodal inputs. The benefits of pre-training MuJo on FiMAD, especially when limited data are available for the target task, are analyzed. Furthermore, we analyze the zero-shot performance of our method and compare unimodal versus multimodal training on all four datasets.
@@ 2 Related Work
Datasets for activity recognition
In recent years, several datasets have been developed to advance HAR for sports activities, supporting various applications and enabling robust evaluations across multiple modalities, including images (Verma et al., 2020), videos (Shao et al., 2020), and wearable sensors (Malekzadeh et al., 2018). Furthermore, cross-modal datasets, such as video-pose (Fieraru et al., 2021) and video-sensor datasets (Strömbäck et al., 2020; Kwon et al., 2021b), have been developed to support transfer learning in HAR (Thukral et al., 2023; Kamboj and Do, 2024), allowing knowledge transfer between different data modalities and enhancing model adaptability. However, these datasets are dependent on human-annotated activity labels, a process that is both costly and time-consuming, posing challenges for large-scale data collection and scalability. For example, in (Kwon et al., 2021b), virtual IMU data was simulated from YouTube videos and combined with real data to improve classification performance. However, this approach required collecting and labeling videos for target activities as well as careful calibration of the generated simulated sensor data. Therefore, we propose a novel data curation method that utilizes YouTube chapters to collect videos annotated with corresponding activity classes, while also expanding the range of modalities (explained in Section 3). This method facilitates the collection of multimodal data encompassing sensor, pose, video, and text features to enable learning of a richer vector space.
Unimodal and multimodal activity recognition
@@
Unimodal activity recognition has been extensively studied in HAR. Various approaches have been proposed to address this task and can be broadly categorized into two groups: handcrafted feature-based and deep learning-based methods. Handcrafted feature-based methods rely on extracting low-level features from the input data and designing a classifier to recognize activities. For instance, in accelerometer-based activity recognition (Kwapisz et al., 2011), features such as the mean, standard deviation, and energy of the acceleration signal are often used to capture the characteristics of different activities. Popular classifiers used in this approach include decision trees, support vector machines, and k-nearest neighbors. Deep learning-based methods, on the other hand, use neural networks to learn features and recognize activities from raw input data. Convolutional neural networks (CNNs) and recurrent neural networks (RNNs) are two popular types of neural networks used in this approach. CNNs are commonly used for image-based activity recognition (Strömbäck et al., 2020), being combined with RNNs in some cases (Ordóñez and Roggen, 2016).
@@
Despite the success of unimodal activity recognition, it is often limited because a single modality may only capture a part of the essential information about the activity, leading to lower recognition accuracy. Therefore, multimodal activity recognition, which combines information from multiple sources, such as IMU data, audio, and video (Mekruksavanich and Jitpattanakul, 2022; Ijaz et al., 2022), has gained increasing attention in recent years. These modalities can be combined using fusion methods such as early or late fusion (Duhme et al., 2022). Some studies have also focused on domain adaptation, where the goal is to recognize activities in a new environment using data from a different environment (Plananamente et al., 2022). In such cases, transfer learning techniques can adapt the model to the new environment using a small amount of labeled data. Overall, multimodal activity recognition has shown promising results, with improved accuracy compared to unimodal approaches. However, challenges in data fusion, feature extraction, and model selection still need to be addressed to improve the performance of multimodal activity recognition systems (Islam et al., 2022).
Representation learning with proxy tasks
@@
Different modalities, e.g., natural language, visual inputs, or sensor signals, often complement a common concept. A key challenge of multimodal feature representation is exploring efficient methods for multimodal fusion with the given modalities to avoid missing modal features. We can differentiate the representation learning methods in two aspects: learning objectives and fusing the learned features in a joint space. In learning objectives, various approaches have been proposed to learn multimodal representations by formulating proxy tasks such as reconstruction, completion (Seo et al., 2021; Sun et al., 2019), matching/alignment (Li et al., 2020; Miech et al., 2020), and ordering tasks (Li et al., 2020). Generating one feature vector with multimodal inputs for a common concept is a crucial problem of multimodal representation learning. Simple methods, such as concatenating learned unimodal feature vectors, have been extensively explored to enhance the performance of downstream tasks (Hu et al., 2018; Strömbäck et al., 2020). These methods involve combining feature vectors from a single modality, such as text or image, to create more informative representations that can be used for various tasks. More advanced methods have also been proposed, including a tensor fusion network (Zadeh et al., 2017) and a transformer-based co-attention mechanism for efficient cross-modality learning (Lu et al., 2019).
@@
Proxy tasks were also developed for the field of HAR with wearable devices including IMUs, where unlabeled IMU data are used to pre-train better representations. Those approaches include Multitask (Saeed et al., 2019), Auto-encoder (Haresamudram et al., 2019) and simCLR (Tang et al., 2021). Multi-task (Saeed et al., 2019) proposes auxiliary tasks of binary classification for extracting useful features for the downstream task. A multitask temporal CNN is trained to classify whether the input signal was transformed or not for each of a set of transformations such as scaling, rotating the signal, etc. The authors utilize a large unlabeled accelerometer dataset for self-supervised learning and transfer the learned knowledge to a small, labeled activity recognition dataset. Auto-encoder (Haresamudram et al., 2019) is the application of auto-encoders, where an encoder network reduces the dimensionality of the data, while a decoder brings the representation back to its original dimensions, with the whole system learning to reconstruct the original signal. simCLR (Tang et al., 2021) learns general representations from a dataset by training a model to match different view instances created from the same signal by contrasting them. The positive pair is obtained by the transformation of the input window, while the negative pair is formed by the remaining input windows of the same batch. Then, the contrastive loss pulls the positive pair together, while pulling the negative pair far away from each other in the representation space.
Representation via contrastive learning
@@
Contrastive learning is an efficient self-supervised framework applied across multiple domains, which learns similar/dissimilar representations from data that are organized into similar and dissimilar pairs. Recently, Radford et al. presented a novel approach, known as CLIP (Radford et al., 2021), to jointly represent images and text descriptions by leveraging contrastive loss with the similarity of image-text pairs. CLIP learns two feature spaces with two encoders, i.e., image and text encoders, then projects the learned features into a shared latent space with contrastive loss. Contrastive learning has also been extended to other modalities. For example, video-text data have attracted interest in studies focused on the shared space of video-text features (Xu et al., 2021; Xue et al., 2022) and text-to-video generation (Singer et al., 2022). MotionCLIP (Tevet et al., 2022) trained an encoder to find the appropriate embedding of an input sequence in the CLIP space and a decoder to generate the most fitting motion for a given CLIP space. Following this, IMU2CLIP (Moon et al., 2022) was proposed to learn a joint space that aligns IMU motion sensor recordings with video and text by projecting them into a shared representation space. Studies in action recognition (Li et al., 2020; Wang et al., 2021; Sun et al., 2022; Morshed et al., 2023) have also significantly enhanced the field of fitness activity recognition, despite traditionally focusing on a broad spectrum of actions such as “cutting the potato” and “cleaning the drawer”. Certain investigations in action recognition also aim to acquire knowledge of the feature space through contrastive learning, as exemplified by initiatives such as HERO (Li et al., 2020) and ActionCLIP (Wang et al., 2021).
@@
The achievements discussed in the aforementioned studies have shown that contrastive learning has a significant impact on extracting a joint space with similar and dissimilar pairs. Consequently, the contrastive learning methodology employed in our study aims to acquire knowledge of the feature space pre-training to fitness activities. Our multimodal contrastive learning includes video, IMU sensors, poses, and text, within a multitask training manner in an end-to-end training paradigm similar to ImageBind (Girdhar et al., 2023). Recent works in multimodal pre-training also make use of simulated sensor data from video sources (Fortes Rey et al., 2024), but focus only on sensors placed on the wrist and do not include video features. Moreover, they do not investigate pre-training on a large amount of YouTube fitness data, relying instead on sign language videos and motion capture datasets.